A shopkeeper buys a stove from a manufacturer for $860. He sells it for $ 650. Calculate his loss percentage.
Answer by Theo(13342)

the cost is 860.
the selling price is 650.
the difference between the selling price and the cost is 210.
since the selling price is less than the cost, the difference is a loss.
the loss percent is the loss divided by the cost = 210/860 = .2441860465 * 100 = 24.42% rounded to 2 decimal places.

another way to calculate this is:

profit = (selling price / cost) minus 1.
if the profit is negative, then make it positive and call it a loss.
650 / 860 = .7558139535 minus 1 = -.2441860465 * 100 = -24.42%.
since the profit is negative, then it's a loss of 24.42%.